# Bandit Learning in Housing Markets

### v1

@@ Abstract
The housing market, also known as one-sided matching market, is a classic exchange economy model where each agent on the demand side initially owns an indivisible good (a house) and has a personal preference over all goods. The goal is to find a core-stable allocation that exhausts all mutually beneficial exchanges among subgroups of agents. While this model has been extensively studied in economics and computer science due to its broad applications, little attention has been paid to settings where preferences are unknown and must be learned through repeated interactions. In this paper, we propose a statistical learning model within the multi-player multi-armed bandit framework, where players (agents) learn their preferences over arms (goods) from stochastic rewards. We introduce the notion of core regret for each player as the market objective. We study both centralized and decentralized approaches, proving $\mathcal{O}(N\log T/\Delta^{2})$ upper bounds on regret, where $N$ is the number of players, $T$ is the time horizon and $\Delta$ is the minimum preference gap among players. For the decentralized setting, we also establish a matching lower bound, demonstrating that our algorithm is order-optimal.
@@ 1 Introduction
The housing market, also known as one-sided matching market, represents a foundational economic institution where agents exchange property rights to achieve more desirable arrangements. Classical research in this domain traces back to the seminal work of shapley1974cores, and the model has been extensively studied in the literature (roth2023online) due to its wide range of applications such as student housing assignments (sonmez2010house), school choice (abdulkadirouglu2003school) and kidney exchange (roth2004kidney). In a housing market, each agent starts with an indivisible good – typically representing a house – and has their own personal preferences over all available goods in the market. Unlike traditional markets with buyers and sellers, here, agents trade their initial endowments based on their preferences, seeking to obtain a house they value more, and the houses must be traded without monetary transactions. The key challenge is to find a stable and fair allocation where no agent can improve their outcome by trading with someone else. The Top Trading Cycle (TTC) algorithm introduced by shapley1974cores is a celebrated mechanism that ensures efficient, strategy-proof and Pareto optimal outcomes (roth1977weak). Subsequent research has expanded this framework to accommodate different settings and provide various solutions to the problem (hylland1979efficient; echenique2021constrained; garg2024one). Despite these advancements, the integration of adaptive learning mechanisms into housing market models remains an underexplored area, particularly in online environments where participants must navigate uncertainty in preferences.
@@
The multi-armed bandit (MAB) framework models how a player learns in an unknown environment with limited feedback (auer2002finite; lattimore2020bandit). In the basic setup, a player faces $K$ arms, each with an initially unknown reward distribution. Upon selecting an arm, the player observes a stochastic reward and updates their belief about the arm’s preference. The goal is to maximize cumulative expected rewards, or equivalently, minimize cumulative expected regret – the difference between rewards from the optimal arm and the player’s chosen arms over time. Classical strategies for balancing exploration (learning arm preferences) and exploitation (leveraging known rewards), such as explore-then-commit (ETC) (garivier2016explore), upper confidence bound (UCB) (auer2002finite) and Thompson sampling (TS) (thompson1933likelihood), achieve sublinear regret, ensuring asymptotic optimality.
@@
To formalize these challenges, we initiate bandit learning in housing markets in this paper, abstract the market as a multiplayer bandit problem. Here, players and arms correspond to agents and houses, each player has heterogeneous and unknown preferences over the arms, and each player is naturally associated with an arm, which represents their initially endowed house. When being matched to an arm, the player could learn the corresponding preference through the stochastic reward, but every arm could be matched to at most one player every time, when more than one player try to pull the same arm, all of them would get collision and receive no rewards. Taking the outcome from the TTC algorithm as a natural and desirable solution for the market, denoted as the core matching, our objective is to minimize the regret defined as the cumulative reward difference between the arm from the core matching and the player’s selected arm. We develop matching and learning algorithms that can provably attain the core of the market in this setting. Our contributions are as follows:
@@
  • When the horizon $T$ is unknown, we provide a UCB-type algorithm in the centralized market setting, which is adaptive and anytime. We prove that centralized UCB achieves ${\mathcal{O}}(N\log T/\Delta_{\min}^{2})$ problem-dependent upper bounds on the regret for every player.
@@ 2 Related Work
The multi-player bandit problem involves multiple decentralized players interacting with a shared multi-armed bandit environment. When players pull the same arm simultaneously, a collision occurs, resulting in a loss. In settings where arm means vary across players, the benchmark is typically the maximum weight matching, and the system regret – defined as the cumulative reward loss summed over all players – measures performances. tibrewal2019distributed proposed an ETC type algorithm where players exploit the best-estimated matching, mehrabian2020practical combines forced collisions for implicit communication with matching eliminations, shi2021heterogeneous adapted the CUCB algorithm (chen2013combinatorial) to this setting. For a comprehensive survey, see boursier2024survey.
While our model adopts the same reward and collision structure as multiplayer bandits with heterogeneous arm means, our benchmark diverges: we use the core of the housing market – a game-theoretic solution concept distinct from maximum weight matching – and define individual regret for each player rather than aggregate system regret. This shift ensures fairness but introduces new algorithmic challenges.
@@
Bandit problems in matching markets were first formalized by das2005two, with subsequent works (liu2020competing; liu2021bandit; sankararaman2021dominate; kong2023player) exploring this model to achieve stable matching (gale1962college). In these two-sided markets, players (with unknown utilities) and arms (with known preferences) interact – when multiple players pull the same arm, only the top-ranked player by the arm’s preference gets matched while others face collisions, with individual regrets defined accordingly. Since multiple stable matchings may exist, regrets are typically measured against either player-optimal or player-pessimal stable matchings.
In contrast, we study housing markets (one-sided matching) where only players have preferences over arms. This creates a distinct collision structure: when multiple players propose to the same arm, no one receives a reward. Moreover, the core matching in housing markets is unique, ensuring our regret definition is unambiguous. These differences from two-sided markets introduce novel algorithmic challenges.
@@ 3 Problem Setting
In this section, we formally introduce the problem setting of bandit learning in housing markets.
@@
The core is a fundamental solution concept in housing markets (shapley1974cores), which is the set of feasible allocations where no coalition of agents can benefit by breaking away from the grand coalition. Formally,
@@
In housing markets with strict preferences, the core is always nonempty, and consists of a unique matching (roth1977weak), which we refer to as the core matching and denote by $\mu^{*}$. Our objective is to learn $\mu^{*}$ while minimizing the core regret for each player $p_{i}\in{\mathcal{N}}$. This regret is defined as the cumulative difference, over $T$ rounds, between the expected rewards from being matched to $\mu^{*}(p_{i})$ and the rewards actually obtained by $p_{i}$:
@@
In the offline setting where all players know their exact preferences, the Top-Trading-Cycle (TTC) algorithm (shapley1974cores) efficiently computes the unique core allocation of the housing market – a stable assignment where no coalition of players can improve their outcomes through mutual exchanges. The centralized TTC proceeds iteratively. First, each player identifies her most preferred available arm. Second, a directed graph is formed where players point to owners of their top choices. Third, at least one cycle (including possible self-cycles) is identified and implemented, with involved players exiting the market. This process terminates within $N$ steps, guaranteed to produce a core allocation.
@@ 4 Decentralized Algorithm
In this section, we present an Explor-then-commit (ETC) algorithm for decentralized housing markets, operating in two phases. First, players explore arms in a round-robin fashion to estimate their preference rankings (Line 3-26), ensuring reliable ordinal estimates for matching. Next, players use these estimates to identify their assigned arms in the core matching via a decentralized adaptation of the You Request My House, I Get Your Turn (YRMH-IGYT) mechanism (abdulkadirouglu1999house) (Line 27-52). Once matched, players commit exclusively to these arms in all future rounds.
Algorithm 1 Decentralized Bandit Learning in Housing Markets (from view of player $p_{i}$)
1:player set $\mathcal{P}$} and the corresponding arm set $\mathcal{A}$}, horizon $T$.
@@
5:  $P_{\ell}=$ False // whether the preference has been estimated well
@@
11:  if $\exists\sigma$ such that $\text{LCB}_{i,\sigma_{k}}>\text{UCB}_{i,\sigma_{k+1}}$ for any $k\in[N]$ then
12:    $P_{\ell}=$ True and $\sigma_{i}=\sigma$.
@@
23:     $t_{1}=\sum_{\ell^{\prime}=1}^{\ell}(2^{\ell^{\prime}}+N)$ //The round when phase 1 ends.
@@
When the confidence intervals of two arms $a_{j},a_{j^{\prime}}$ become disjoint, that is, wehn $\text{LCB}_{i,j}>\text{UCB}_{i,j^{\prime}}$ or vice versa, player $p_{i}$ can establish its strict preference ordering between them. Once $p_{i}$ successfully determines a complete preference ranking over all $N$ arms (Line 11), it sets the flag $P_{\ell}$ to True for the current sub-phase and records this permutation as its estimated preference ranking $\sigma_{i}$.
Communication rounds enable players verify if all participants have accurately estimated their preference rankings. In the $i$-th communication round of every sub-phase, player $p_{i}$’s endowed arm serves as a broadcast channel. A player $p_{j}$ proposes to $p_{i}$’s arm if and only if it has an accurate ranking, i.e., $P_{\ell}=$ True (Line 17); otherwise, it abstains (Line 19). While this may cause collisions, the goal is not to resolve them but to convey status information. If $p_{i}$ observes proposals from all $N$ players including itself (Line 21), it infers universal estimation success and triggers the transition to the second phase using its estimated ranking $\sigma_{i}$ (Line 22).
The second phase (Line 27-52) adapts the YRMH-IGYT algorithm (abdulkadirouglu1999house) to our decentralized setting for players to progressively identify their core allocations. The process unfolds in successive sub-phases. In each, the unassigned player with the smallest index initiates a proposal chain by proposing to their most preferred remaining arm (Lin 37-38). The recipient of a proposal then proposes to their own top choice among available arms, and this sequence continues. A top trading cycle is completed when a proposal is received by a player who is already part of the current chain (Line 44). The player who closes the cycle – the first to receive a repeat proposal – then marks their endowed arm as allocated. This allocation status propagates backward through the chain. All players involved in the cycle permanently fix their allocations and remove their own endowed arms from the market (Line 45). The process repeats among the remaining players until all arms are allocated, thus achieving a stable core matching in a fully decentralized manner.
@@ 4.1 Theoretical Analysis
The regret bound consists of three key components. First, the cumulative regret from exploration rounds in phase 1. Second, the regret from communication rounds in phase 1. The final term combines two distinct elements: (1) the regret during phase 2, where the YRMH-IGYT mechanism guarantees convergence to the core matching within at most $N$ epoches (with each epoch requirting $\leq N$ rounds to identify a top trading cycle), and (2) the regret contribution from rare concentration failure events.
@@
$\leq$ $\mathbb{E}\left[\sum_{t=1}^{T}\mathds{1}\left\{\mu_{t}(i)\neq\mu^{*}(p_{i})\right\}\middle|\urcorner{\mathcal{F}}\right]\cdot\Delta_{i,\max}+{\mathbb{P}}({\mathcal{F}})\cdot T\cdot\Delta_{i,\max}$
@@
$\leq$ $\left(\frac{192N\log T}{\Delta_{\min}^{2}}+N\log\left(\frac{192N\log T}{\Delta_{\min}^{2}}\right)\right)\cdot\Delta_{i,\max}+3N^{2}\Delta_{i,\max},$ (8)
@@
The following the technical lemmas underlying our analysis. Proofs are provided in Appendix A. We begin with Lemma 1, which bounds the probability of erroneous preference estimation.
@@
which implies that $\sum_{\ell^{\prime}=1}^{\ell_{\max}}2^{\ell^{\prime}}\leq\frac{192N\log T}{\Delta_{\min}^{2}}$ and $\ell_{\max}=\log\left(\frac{192N\log T}{\Delta^{2}}\right)$ since the sub-phase length grows exponentially. And all players will enter in phase 2 simultaneously at the end of sub-phase $\ell_{\max}$.
@@ 5 Regret Lower Bound
We establish a regret lower bound by adapting the method of burnetas1996optimal to our multi-player setting, demonstrating the tightness of the regret upper bound of Algorithm 1. Let $Reg(T;\bm{\nu},\pi)$ be the cumulative expected regret of policy $\pi$ over all players and time $T$, for an instance with arm distributions $\bm{\nu}=\left\{\nu_{ij}:i,j\in[N]\right\}$. Denote by $\mathcal{P}$} the set of all probability distributions with support in $[0,1]$. For any $\bm{\nu}\in{\mathcal{P}}$, define $D_{\inf}(\bm{\nu},x,{\mathcal{P}})=\inf_{\bm{\nu}^{\prime}\in{\mathcal{P}}}\left\{D(\bm{\nu},\bm{\nu}^{\prime}):\rho(\bm{\nu}^{\prime})>x\right\}$, where $\rho:{\mathcal{P}}\rightarrow{\mathbb{R}}$ maps a distribution to its mean and $D(\cdot,\cdot)$ is the KL divergence.
@@
This ensures the policy is not overly tuned to the current instance at the expense of performance in others, ensuring robust performance across different problem instances (burnetas1996optimal; lattimore2020bandit).
@@
For any player $p_{i}$, $i\in[N]$, under any decentralized universally consistent algorithm $\pi$ on an STTCB instance $\bm{\nu}$ satisfies
$\begin{split}&\liminf_{T\to\infty}\frac{Reg_{i}(T;\bm{\nu},\pi)}{\log T}\\ \geq&\max\Biggl\{\sum_{i^{\prime}\neq i}\frac{\Delta_{\min}^{(i)}}{D_{\inf}(\bm{\nu}_{i^{\prime},\mu^{*}(p_{i})},{\bm{U}}(i^{\prime},\mu^{*}(p_{i^{\prime}})),{\mathcal{P}})},\sum_{a_{j}\neq\mu^{*}(p_{i})}\frac{\Delta_{j}^{(i)}}{D_{\inf}(\bm{\nu}_{i,j},{\bm{U}}(i,\mu^{*}(p_{i})),{\mathcal{P}})}\Biggr\}.\end{split}$ (10)
Proof Idea of Theorem 2.
@@
The detailed proof of Theorem 2 is deferred to Appendix B. The following corollary demonstrates that the $N\log T/\Delta_{\min}^{2}$ dependence in Theorem 1 cannot be improved, thus establishing $\Theta(N\log T/\Delta_{\min}^{2})$ as the minimax regret rate for all players.

### v2

@@ Abstract
The housing market, also known as one-sided matching market, is a classic exchange economy model where each agent on the demand side initially owns an indivisible good (a house) and has a personal preference over all goods. The goal is to find a core-stable allocation that exhausts all mutually beneficial exchanges among subgroups of agents. While this model has been extensively studied in economics and computer science due to its broad applications, little attention has been paid to settings where preferences are unknown and must be learned through repeated interactions. In this paper, we propose a statistical learning model within the multi-player multi-armed bandit framework, where players (agents) learn their preferences over arms (goods) from stochastic rewards. We introduce the notion of core regret for each player as the market objective. We study both centralized and decentralized approaches, proving $\mathcal{O}(\log T/\Delta^{2})$ upper bounds on regret, where $T$ is the time horizon and $\Delta$ is the minimum preference gap among players. For the decentralized setting, we also establish a matching lower bound, demonstrating that our algorithm is order-optimal.
@@ 1 Introduction
The housing market, also known as one-sided matching market, represents a foundational economic institution where agents exchange property rights to achieve more desirable arrangements. Classical research in this domain traces back to the seminal work of Shapley and Scarf (1974), and the model has been extensively studied in the literature (Roth 2023) due to its wide range of applications such as student housing assignments (Sönmez and Ünver 2010), school choice (Abdulkadiroğlu and Sönmez 2003) and kidney exchange (Roth et al. 2004). In a housing market, each agent starts with an indivisible good – typically representing a house – and has their own personal preferences over all available goods in the market. Unlike traditional markets with buyers and sellers, here, agents trade their initial endowments based on their preferences, seeking to obtain a house they value more, and the houses must be traded without monetary transactions. The key challenge is to find a stable and fair allocation where no agent can improve their outcome by trading with someone else. The Top Trading Cycle (TTC) algorithm introduced by Shapley and Scarf (1974) is a celebrated mechanism that ensures efficient, strategy-proof and Pareto optimal outcomes (Roth and Postlewaite 1977). Subsequent research has expanded this framework to accommodate different settings and provide various solutions to the problem (Hylland and Zeckhauser 1979; Echenique et al. 2021; Garg et al. 2024). Despite these advancements, the integration of adaptive learning mechanisms into housing market models remains an underexplored area, particularly in online environments where participants must navigate uncertainty in preferences.
@@
The multi-armed bandit (MAB) framework models how a player learns in an unknown environment with limited feedback (Auer et al. 2002a; Lattimore and Szepesvári 2020). In the basic setup, a player faces $K$ arms, each with an initially unknown reward distribution. Upon selecting an arm, the player observes a stochastic reward and updates their belief about the arm’s preference. The goal is to maximize cumulative expected rewards, or equivalently, minimize cumulative expected regret – the difference between rewards from the optimal arm and the player’s chosen arms over time. Classical strategies for balancing exploration (learning arm preferences) and exploitation (leveraging known rewards), such as explore-then-commit (ETC) (Garivier et al. 2016), upper confidence bound (UCB) (Auer et al. 2002a) and Thompson sampling (TS) (Thompson 1933), achieve sublinear regret, ensuring asymptotic optimality.
@@
To formalize these challenges, we initiate bandit learning in housing markets, abstract the market as a multiplayer bandit problem. Here, players and arms correspond to agents and houses, each player has heterogeneous and unknown preferences over the arms, and each player is associated with an arm, which denotes their initially endowed house. When being matched to an arm, the player could learn the corresponding preference through the stochastic reward, but every arm could be matched to at most one player every time, when more than one player try to pull the same arm, all of them would get collision and receive no rewards. Taking the outcome from the TTC algorithm as a natural and desirable solution for the market, denoted as the core matching, our objective is to minimize the regret defined as the cumulative reward difference between the arm from the core matching and the player’s selected arm. We develop matching and learning algorithms that can provably attain the core of the market in this setting. Our contributions are as follows:
@@
  • When the horizon $T$ is unknown, we provide a UCB-type algorithm in the centralized market setting, which is adaptive and anytime. We prove that centralized UCB achieves ${\mathcal{O}}(N^{2}\log T/\Delta_{\min}^{2})$ problem-dependent upper bounds on the regret for every player.
@@ 2 Related Work
The multi-player bandit problem involves multiple decentralized players interacting with a shared multi-armed bandit environment. When players pull the same arm simultaneously, a collision occurs, resulting in a loss. In settings where arm means vary across players, the benchmark is typically the maximum weight matching, and the system regret – defined as the cumulative reward loss summed over all players – measures performances. Tibrewal et al. (2019) proposed an ETC type algorithm where players exploit the best-estimated matching, Mehrabian et al. (2020) combined forced collisions for implicit communication with matching eliminations, Shi et al. (2021) adapted the CUCB algorithm (Chen et al. 2013) to this setting. For a comprehensive survey, see Boursier and Perchet (2024).
While our model shares the reward and collision structure of heterogeneous multiplayer bandits, our benchmark differs fundamentally: we use the core of the housing market —- a game-theoretic solution distinct from maximum weight matching – and define individual regret per player rather than aggregate system regret. This ensures fairness but introduces distinct algorithmic challenges.
@@
Bandit problems in matching markets were first formalized by Das and Kamenica (2005), with subsequent works (Liu et al. 2020, 2021; Sankararaman et al. 2021; Kong and Li 2023; Lin et al. 2025) exploring this model to achieve stable matching (Gale and Shapley 1962). In these two-sided markets, players (with unknown utilities) and arms (with known preferences) interact – when multiple players pull the same arm, only the top-ranked player by the arm’s preference gets matched while others face collisions, with individual regrets defined accordingly. Since multiple stable matchings may exist, regrets are typically measured against either player-optimal or player-pessimal stable matchings.
In contrast, we study housing markets (one-sided matching) where only players have preferences over arms. This creates a distinct collision structure: when multiple players propose to the same arm, all receive zero reward. Moreover, the core matching in a housing markets is unique, ensuring an unambiguous regret definition. These differences from two-sided markets introduce novel algorithmic challenges.
@@ 3 Problem Setting
The core is a fundamental solution concept in housing markets (Shapley and Scarf 1974), which is the set of feasible allocations where no coalition of agents can benefit by breaking away from the grand coalition. Formally,
@@
In housing markets with strict preferences, the core is always nonempty, and consists of a unique matching (Roth and Postlewaite 1977), which we refer to as the core matching and denote by $\mu^{*}$. Our objective is to learn $\mu^{*}$ while minimizing the core regret for each player $p_{i}\in{\mathcal{N}}$. This regret is defined as the cumulative difference, over $T$ rounds, between the expected rewards from being matched to $\mu^{*}(p_{i})$ and the rewards actually obtained by $p_{i}$:
@@
In the offline setting where all players know their exact preferences, the Top-Trading-Cycle (TTC) algorithm (Shapley and Scarf 1974) efficiently computes the unique core allocation of the housing market – a stable assignment where no coalition of players can improve their outcomes through mutual exchanges. The centralized TTC proceeds iteratively. First, each player identifies her most preferred available arm. Second, a directed graph is formed where players point to owners of their top choices. Third, at least one cycle (including possible self-cycles) is identified and implemented, with involved players exiting the market. This process terminates within $N$ steps, guaranteed to produce a core allocation.
@@ 4 Decentralized Algorithm
In this section, we present an Explore-then-Commit (ETC) algorithm for decentralized housing markets, operating in two phases. First, players explore arms in a round-robin fashion to estimate their preference rankings (Line 3-26), ensuring reliable ordinal estimates for matching. Next, players use these estimates to identify their assigned arms in the core matching via a decentralized adaptation of the You Request My House, I Get Your Turn (YRMH-IGYT) mechanism (Abdulkadiroğlu and Sönmez 1999) (Line 27-52). Once matched, players commit exclusively to these arms in all future rounds.
Algorithm 1 Decentralized Explore-then-YRMH-IGYT (from view of player $p_{i}$)
1:player set $\mathcal{P}$} and arm set $\mathcal{A}$}, horizon $T$.
@@
5:  $P_{\ell}^{(i)}=$ False //whether the preference is well-estimated
@@
11:  if $\exists\sigma$ such that $\text{LCB}_{i,\sigma_{k}}>\text{UCB}_{i,\sigma_{k+1}}$, $\forall k\in[N]$ then
12:   $P_{\ell}^{(i)}=$ True and $\sigma_{i}=\sigma$.
@@
23:     $t_{1}=\sum_{\ell^{\prime}=1}^{\ell}(2^{\ell^{\prime}}+N)$ //Phase 1 ends.
@@
When the confidence intervals of two arms $a_{j},a_{j^{\prime}}$ become disjoint, that is, when $\text{LCB}_{i,j}>\text{UCB}_{i,j^{\prime}}$ or vice versa, player $p_{i}$ can establish its strict preference ordering between them. Once $p_{i}$ successfully determines a complete preference ranking over all $N$ arms (Line 11), it sets the flag $P_{\ell}^{(i)}$ to True for the current sub-phase and records this permutation as its estimated preference ranking $\sigma_{i}$.
Communication rounds enable players verify if all participants have accurately estimated their preference rankings. In the $i$-th communication round of every sub-phase, player $p_{i}$’s endowed arm serves as a broadcast channel. A player $p_{j}$ proposes to $p_{i}$’s arm if and only if it has an accurate ranking (Line 17); otherwise, it abstains (Line 19). While this may cause collisions, the goal is not to resolve them but to convey status information. If $p_{i}$ observes proposals from all $N$ players including itself (Line 21), it infers universal estimation success and triggers the transition to the second phase using its estimated ranking $\sigma_{i}$ (Line 22).
The second phase (Line 27-52) adapts the YRMH-IGYT algorithm (Abdulkadiroğlu and Sönmez 1999) to our decentralized setting for players to progressively identify their core allocations. The process unfolds in successive sub-phases. In each, the unassigned player with the smallest index initiates a proposal chain by proposing to their most preferred remaining arm (Lin 37-38). The recipient of a proposal then proposes to their own top choice among available arms, and this sequence continues. A top trading cycle is completed when a proposal is received by a player who is already part of the current chain (Line 44). The player who closes the cycle – the first to receive a repeat proposal – then marks their endowed arm as allocated. This allocation status propagates backward through the chain. All players involved in the cycle permanently fix their allocations and remove their own endowed arms from the market (Line 45). The process repeats among the remaining players until all arms are allocated, thus achieving a stable core matching in a fully decentralized manner.
@@
The regret bound consists of three key components. First, the cumulative regret from exploration rounds in phase 1. Second, the regret from communication rounds in phase 1. The final term combines two distinct elements: (1) the regret during phase 2, where the YRMH-IGYT mechanism guarantees convergence to the core matching within at most $N$ epochs (with each epoch requiring $\leq N$ rounds to identify a top trading cycle), and (2) the regret contribution from rare concentration failure events.
@@
$\leq$ $\mathbb{E}\left[\sum_{t=1}^{T}\mathds{1}\left\{\mu_{t}(i)\neq\mu^{*}(p_{i})\right\}\middle|\urcorner{\mathcal{F}}\right]\cdot\Delta_{i,\max}$
$+{\mathbb{P}}({\mathcal{F}})\cdot T\cdot\Delta_{i,\max}$
@@
$\leq$ $\left(\frac{192N\log T}{\Delta_{\min}^{2}}+N\log\left(\frac{192N\log T}{\Delta_{\min}^{2}}\right)\right)\cdot\Delta_{i,\max}$
$+3N^{2}\Delta_{i,\max},$ (8)
@@
The following technical lemmas underlying our analysis. Proofs are provided in Appendix A. We begin with Lemma 1, which bounds the probability of erroneous preference estimation.
@@
which implies that $\sum_{\ell^{\prime}=1}^{\ell_{\max}}2^{\ell^{\prime}}\leq\frac{192N\log T}{\Delta_{\min}^{2}}$ and $\ell_{\max}=\log\left(\frac{192N\log T}{\Delta_{\min}^{2}}\right)$ since the sub-phase length grows exponentially. And all players will enter in phase 2 simultaneously at the end of sub-phase $\ell_{\max}$.
@@ 5 Regret Lower Bound
We establish a regret lower bound by adapting the method of Auer et al. (2002b) to our multi-player setting, demonstrating the tightness of the regret upper bound of Algorithm 1. Let $Reg(T;\bm{\nu},\pi)$ be the cumulative expected regret of policy $\pi$ over all players and time $T$, for an instance with arm distributions $\bm{\nu}=\left\{\nu_{ij}:i,j\in[N]\right\}$. Denote by $\mathcal{P}$} the set of all probability distributions with support in $[0,1]$. For any $\bm{\nu}\in{\mathcal{P}}$, define $D_{\inf}(\bm{\nu},x,{\mathcal{P}})=\inf_{\bm{\nu}^{\prime}\in{\mathcal{P}}}\left\{D(\bm{\nu},\bm{\nu}^{\prime}):\rho(\bm{\nu}^{\prime})>x\right\}$, where $\rho:{\mathcal{P}}\rightarrow{\mathbb{R}}$ maps a distribution to its mean and $D(\cdot,\cdot)$ is the KL divergence.
@@
This ensures the policy is not overly tuned to the current instance at the expense of performance in others, ensuring robust performance across different problem instances (Auer et al. 2002b; Lattimore and Szepesvári 2020).
@@
$Reg_{i}(T;\bm{\nu},\pi)\geq\max\Biggl\{$ $\sum_{i^{\prime}\neq i}\Delta_{\min}^{(i)}\mathbb{E}_{\bm{\nu},\pi}\left[N_{\mu^{*}(p_{i})}^{(i^{\prime})}(T)\right],$
$\sum_{a_{j}\neq\mu^{*}(p_{i})}\Delta_{j}^{(i)}\mathbb{E}_{\bm{\nu},\pi}\left[N_{j}^{(i)}(T)\right]\Biggr\}.$
@@
For any player $p_{i}$, $i\in[N]$, and under any decentralized uniformly consistent algorithm $\pi$, the performance on an STTCB instance $\bm{\nu}$ satisfies
$\begin{split}&\liminf_{T\to\infty}\frac{Reg_{i}(T;\bm{\nu},\pi)}{\log T}\\ \geq\max\Biggl\{&\sum_{i^{\prime}\neq i}\frac{\Delta_{\min}^{(i)}}{D_{\inf}(\bm{\nu}_{i^{\prime},\mu^{*}(p_{i})},{\bm{U}}(i^{\prime},\mu^{*}(p_{i^{\prime}})),{\mathcal{P}})},\\ &\sum_{a_{j}\neq\mu^{*}(p_{i})}\frac{\Delta_{j}^{(i)}}{D_{\inf}(\bm{\nu}_{i,j},{\bm{U}}(i,\mu^{*}(p_{i})),{\mathcal{P}})}\Biggr\}.\end{split}$ (10)
Proof Idea.
@@
The detailed proof of Theorem 2 can be found in Appendix B. The following corollary demonstrates that the $N\log T/\Delta_{\min}^{2}$ dependence in Theorem 1 cannot be improved, thus establishing $\Theta(N\log T/\Delta_{\min}^{2})$ as the minimax regret rate for all players.